A 5% semiannual coupon $100 bond maturing in 15 years is callable on any coupon date after the 10th. If called on the 11th through 20th coupon date, the redemption value would be $110. If callable on the 21st through 30th coupon date, redemption would be at par. Find the price that would ensure an investor a minimum yield of 3% per annum compounded semiannually
Answer #1

The price would be calculated using a) assuming the bond is called on 11th coupon date and b) assuming the bond is called on 21st coupon date and c) assuming the bond is called on 30th coupon date

Case a:
=5%*100/3%*(1-1/1.015^11)+110/1.015^11=118.560450117316

Case b:
=5%*100/3%*(1-1/1.015^21)+100/1.015^21=117.900136734071

Case c:
=5%*100/3%*(1-1/1.015^30)+100/1.015^30=124.015838006234

Price to be paid is 117.90